# latexfontFont shape

### v1

@@ 5.4. Büchi  case, or the general case
Proof 5.6.
Intuitively, the idea is to build an automaton that will make a case disjunction over the different even parities $l$ of $\mathcal{A}$, and will ensure that the run never encounters any priority $>l$ after some time. This is a classical way to turn a parity automaton into a Büchi one. We just need to ensure that the non-determinism introduced in this construction preserve $\omega$-explorability.
Let us first describe formally the construction. We define, for $l$ even parity of $\mathcal{A}$, a copy $\mathcal{A}_{l}$ of $\mathcal{A}$ where all transitions of priority $<l$ become non-Büchi transitions, all transitions of priority $l$ become Büchi transitions, and all transitions of priority $>l$ are rerouted towards a rejecting sink state $\bot$. The automaton $\mathcal{A}_{l}$ recognizes the language of words of $\mathcal{L}(\mathcal{A})$ which can be accepted in $\mathcal{A}$ with infinitely many priorities $l$ and never encounter any priority $>l$.
@@
[…]
$q_{0}$
$p^{\prime}$
$q^{\prime}$
$\mathcal{A}^{\prime}$
$p_{2}$
$q_{2}$
$\mathcal{A}_{2}$
$p_{4}$
$q_{4}$
$\mathcal{A}_{4}$
$p_{6}$
$q_{6}$
$\mathcal{A}_{6}$
$\bot$
Figure 8. Illustration of the construction of $\mathcal{B}$ from $\mathcal{A}$ for the proof of Theorem 33. All transitions associated to a transition $p\overset{a:4}{\longrightarrow}q$ are represented (label $a$ omitted), in bold green for the Büchi transition.
It is clear that $L(\mathcal{B})=L(\mathcal{A})$: an accepting run in $\mathcal{B}$ must jump to some $\mathcal{A}_{l}$ at some point, and from there witness that the word is accepted in $\mathcal{A}$ with priority $l$. Conversely, a $l$-accepting run in $\mathcal{A}$ can be simulated in $\mathcal{B}$ by jumping to the corresponding $\mathcal{A}_{l}$ after the last priority $>l$ is encountered.
@@
    Since some token $i$ accepts in $\sigma$, by seeing infinitely many priority $l$, with no priority $>l$ after some time $j$, the token $t_{i,j,l}$ will accept according to $\sigma^{\prime}$.
@@ Conclusion
Acknowledgments. We thank Milla Valnet for her preliminary work on the explorability question.

### v2

@@ 5.4. Büchi  case, or the general case
Proof 5.8.
Intuitively, the idea is to build an automaton that will make a case disjunction over the different even parities $l$ of $\mathcal{A}$, and will ensure that the run never encounters any priority >l>l after some time. This is a classical way to turn a parity automaton into a Büchi one. We just need to ensure that the non-determinism introduced in this construction preserve $\omega$-explorability.
Let us first describe formally the construction. We define, for $l$ even parity of $\mathcal{A}$, a copy $\mathcal{A}_{l}$ of $\mathcal{A}$ where all transitions of priority <l<l become non-Büchi transitions, all transitions of priority $l$ become Büchi transitions, and all transitions of priority >l>l are rerouted towards a rejecting sink state $\bot$. The automaton $\mathcal{A}_{l}$ recognizes the language of words of $\mathcal{L}(\mathcal{A})$ which can be accepted in $\mathcal{A}$ with infinitely many priorities $l$ and never encounter any priority >l>l.
@@
The automaton $\mathcal{B}$ will simply be the union of $\mathcal{A}^{\prime}$ and of all the $\mathcal{A}_{l}$: the run starts in $\mathcal{A}^{\prime}$, then can non-deterministically jump to any $\mathcal{A}_{l}$ at any time, keeping the local state coherent. The transitions of $\mathcal{B}$ are those of $\mathcal{A}^{\prime},\mathcal{A}_{l}$, plus transition of the form $p^{\prime}\overset{a}{\longrightarrow}q_{l}$ with $p^{\prime}\in\mathcal{A}^{\prime}$ and $q_{l}\in\mathcal{A}_{l}$ for some $l$, corresponding to a transition $p\overset{a}{\longrightarrow}q$ in the original automaton $\mathcal{A}$, regardless of priorities.
It is clear that $L(\mathcal{B})=L(\mathcal{A})$: an accepting run in $\mathcal{B}$ must jump to some $\mathcal{A}_{l}$ at some point, and from there witness that the word is accepted in $\mathcal{A}$ with priority $l$. Conversely, a $l$-accepting run in $\mathcal{A}$ can be simulated in $\mathcal{B}$ by jumping to the corresponding $\mathcal{A}_{l}$ after the last priority >l>l is encountered.
@@
    Since some token $i$ accepts in $\sigma$, by seeing infinitely many priority $l$, with no priority >l>l after some time $j$, the token $t_{i,j,l}$ will accept according to $\sigma^{\prime}$.
@@ Conclusion
Acknowledgments. We thank Milla Valnet for her preliminary work on the explorability question, as well as anonymous reviewers for their helpful feedback.